# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ A.3 Convergence Analysis
Let $H_{j}=\nabla^{2}f(z_{j})+\mu I$. By the fact that $\lambda_{\max}(H_{j}^{-1})=(\lambda_{\min}(H_{j}))^{-1}\leq\frac{1}{\mu-\ell}$ and the definition of $f_{j}$, we have $\|z_{j+1}-z_{j}\|^{2}\geq(\mu-l)(z_{j+1}-z_{j})^{\top}H_{j+1}^{-1}(z_{j+1}-z_{j})=\frac{\mu-l}{\mu^{2}}\|\nabla f(z_{j+1})-\nabla f_{j}(z_{j+1})\|^{2}_{H_{j+1}^{-1}}$. Thus, we have $\|z_{j+1}-z_{j}\|^{2}\geq\frac{\mu-l}{\mu^{2}}(\|\nabla f(z_{j+1})\|_{H_{j+1}^{-1}}-\|\nabla f_{j}(z_{j+1})\|_{H_{j+1}^{-1}})^{2}=\frac{\mu-l}{\mu^{2}}(\lambda_{f_{j+1}}(z_{j+1})-\lambda_{f_{j}}(z_{j+1}))^{2}\geq\frac{\mu-l}{\mu^{2}}(\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\epsilon-\epsilon^{\prime})^{2}=\frac{\mu-l}{4\mu^{2}}\epsilon^{2}$, where the last inequality follows from Theorem A.10. Therefore, we have $\|z_{j+1}-z_{j}\|^{2}\geq\frac{\mu-l}{4\mu^{2}}\epsilon^{2}$ and thus $N\leq\tfrac{8\mu(f(z_{0})-\inf f)}{(\mu-\ell)\epsilon^{2}}$ by Eq. (23).
@@
Let $\Lambda_{1}=C_{\mathrm{CG}}\min\left\{n,\,2Be^{C_{1}\Delta_{f}}\log(\alpha_{\star}^{-1})\right\}$, $\Lambda_{2}=C_{\mathrm{CG}}\min\left\{n,\,2BC_{3}^{6}\log(\alpha_{\star}^{-1})\right\}$, and $\Lambda_{3}=C_{L}(\min\{n,\allowbreak\,\beta^{\prime}\log(n\delta^{\prime-2})\}+\min\{n,\,2Be^{C_{1}\Delta_{f}}\log(n\delta^{\prime-2})\})$. For $j\leq N$, we define $\Delta_{f}^{(j)}=f_{j}(z_{j})-f_{j}(z_{j+1})$ and $K_{1}^{(j)}=\frac{\Delta_{f}^{(j)}}{\omega(R_{3})}$. Moreover, we define $\Delta_{f}^{(N+1)}=0$ and $K_{2}=2\log_{C_{2}^{-1}}\frac{R_{2}}{\kappa\epsilon^{\prime}}$. Theorem A.10 yields that the total number of HVP computations by IPPM is bounded by $\sum_{j=0}^{N+1}(1+K_{1}^{(j)})(1+\Lambda_{1})+(1+K_{2})(1+\Lambda_{2})+\Lambda_{3}\leq\tfrac{1+\Lambda_{1}}{\omega(R_{3})}\sum_{j=0}^{N+1}\Delta_{f}^{(j)}+(1+\Lambda_{2})(N+2)2\log_{C_{2}^{-1}}\frac{R_{2}}{\kappa\epsilon^{\prime}}+(N+2)(2+\Lambda_{1}+\Lambda_{2}+\Lambda_{3})$. The desired result follows from $\sum_{j=0}^{N+1}\Delta_{f}^{(j)}=\sum_{j=0}^{N+1}(f_{j}(z_{j})-f_{j}(z_{j+1}))\allowbreak=\sum_{j=0}^{N+1}(f(z_{j})-f(z_{j+1})-\frac{1}{2}\mu\|z_{j+1}-z_{j}\|^{2})\leq f(z_{0})-\inf f$. ∎

### v2

@@ A.3 Convergence Analysis
Let $H_{j}=\nabla^{2}f(z_{j})+\mu I$. By the fact that $\lambda_{\max}(H_{j}^{-1})=(\lambda_{\min}(H_{j}))^{-1}\leq\frac{1}{\mu-\ell}$ and the definition of $f_{j}$, we have $\|z_{j+1}-z_{j}\|^{2}\geq(\mu-\ell)(z_{j+1}-z_{j})^{\top}H_{j+1}^{-1}(z_{j+1}-z_{j})=\frac{\mu-\ell}{\mu^{2}}\|\nabla f(z_{j+1})-\nabla f_{j}(z_{j+1})\|^{2}_{H_{j+1}^{-1}}$. Thus, we have $\|z_{j+1}-z_{j}\|^{2}\geq\frac{\mu-\ell}{\mu^{2}}(\|\nabla f(z_{j+1})\|_{H_{j+1}^{-1}}-\|\nabla f_{j}(z_{j+1})\|_{H_{j+1}^{-1}})^{2}=\frac{\mu-\ell}{\mu^{2}}(\lambda_{f_{j+1}}(z_{j+1})-\lambda_{f_{j}}(z_{j+1}))^{2}\geq\frac{\mu-\ell}{\mu^{2}}(\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\epsilon-\epsilon^{\prime})^{2}=\frac{\mu-\ell}{4\mu^{2}}\epsilon^{2}$, where the last inequality follows from Theorem A.10. Therefore, we have $\|z_{j+1}-z_{j}\|^{2}\geq\frac{\mu-\ell}{4\mu^{2}}\epsilon^{2}$ and thus $N\leq\tfrac{8\mu(f(z_{0})-\inf f)}{(\mu-\ell)\epsilon^{2}}$ by Eq. (20).
@@
Let $\Lambda_{1}=C_{\mathrm{CG}}\min\left\{n,\,2Be^{C_{1}\Delta_{f}}\log(\alpha_{\star}^{-1})\right\}$, $\Lambda_{2}=C_{\mathrm{CG}}\min\left\{n,\,2BC_{3}^{6}\log(\alpha_{\star}^{-1})\right\}$, and $\Lambda_{3}=C_{L}(\min\{n,\allowbreak\,\beta^{\prime}\log(n\delta^{\prime-2})\}+\min\{n,\,2Be^{C_{1}\Delta_{f}}\log(n\delta^{\prime-2})\})$. For $j\leq N$, we define $\Delta_{f}^{(j)}=f_{j}(z_{j})-f_{j}(z_{j+1})$ and $K_{1}^{(j)}=\frac{\Delta_{f}^{(j)}}{\omega(R_{3})}$. Moreover, we define $\Delta_{f}^{(N+1)}=0$ and $K_{2}=2\log_{C_{2}^{-1}}\frac{R_{2}}{\kappa\epsilon^{\prime}}$. Theorem A.10 yields that the total number of HVP computations by IPPM is bounded by $\sum_{j=0}^{N+1}[(1+K_{1}^{(j)})(1+\Lambda_{1})+(1+K_{2})(1+\Lambda_{2})+\Lambda_{3}]\leq\tfrac{1+\Lambda_{1}}{\omega(R_{3})}\sum_{j=0}^{N+1}\Delta_{f}^{(j)}+(1+\Lambda_{2})(N+2)2\log_{C_{2}^{-1}}\frac{R_{2}}{\kappa\epsilon^{\prime}}+(N+2)(2+\Lambda_{1}+\Lambda_{2}+\Lambda_{3})$. The desired result follows from $\sum_{j=0}^{N+1}\Delta_{f}^{(j)}=\sum_{j=0}^{N+1}(f_{j}(z_{j})-f_{j}(z_{j+1}))\allowbreak=\sum_{j=0}^{N+1}(f(z_{j})-f(z_{j+1})-\frac{1}{2}\mu\|z_{j+1}-z_{j}\|^{2})\leq f(z_{0})-\inf f$. $\square$